# A new web of dualities from Majorana Fermions

### v2

@@ VIII Pairing transition
where $\psi^{2}\equiv\epsilon^{\alpha\beta}\psi_{\alpha}\psi_{\beta}$} with $\epsilon^{12}=-\epsilon^{21}=1$. On the bosonic side, using the identification of the Dirac operator with a monopole operator [Karch:2016sxi], the quadratic operator maps as
@@
$\mathcal{L}_{z}$ $=$ $|D_{a}z|^{2}-|z|^{4}+\frac{1}{4\pi}a\wedge da+\frac{1}{2\pi}A\wedge da$ (39)
$-\big(h\varphi\,\mathcal{M}_{2}+\text{c.c.}\big)-v(\varphi)\,.$
@@
Since $\mathscr{M}_{a}$ is unitary [Turner:2019wnh],
@@
$\mathcal{L}_{\phi}=|D_{a}\vec{\phi}_{1}|^{2}+|D_{a}\vec{\phi}_{2}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})\\ -h\Big[\varphi_{R}\big(\vec{\phi}_{1}^{2}-\vec{\phi}_{2}^{2}\big)-2\varphi_{I}\vec{\phi}_{1}\cdot\vec{\phi}_{2}\Big]\\ -v(\varphi)+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}\,.$ (51)
@@
$\mathcal{L}_{\phi}=|D_{a}\vec{\phi}_{1}|^{2}+|D_{a}\vec{\phi}_{2}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}\\ +N\text{CS}_{g}-h\varphi_{R,0}\big(\vec{\phi}_{1}^{2}-\vec{\phi}_{2}^{2}\big)\,.$ (53)
@@
$\mathcal{L}_{\phi}=|D_{a}\vec{\phi}_{1}|^{2}+|D_{a}\vec{\phi}_{2}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}\\ +N\text{CS}_{g}+2h\varphi_{I,0}\vec{\phi}_{1}\cdot\vec{\phi}_{2}\,.$ (55)
@@
$S_{\phi}=\int d^{3}x\,\Big[\partial_{\mu}\vec{\phi}_{1}\cdot\partial^{\mu}\vec{\phi}_{1}+\partial_{\mu}\vec{\phi}_{2}\cdot\partial^{\mu}\vec{\phi}_{2}\\ +2h\varphi_{I,0}\vec{\phi}_{1}\cdot\vec{\phi}_{2}+\dots\Big]\,.$ (56)
@@
$S_{\phi}$ $=$ $\int\frac{d^{3}p}{(2\pi)^{3}}\begin{pmatrix}\vec{\phi}_{1}^{\mathrm{T}}(p)&\vec{\phi}_{2}^{\mathrm{T}}(p)\end{pmatrix}\begin{bmatrix}-p^{2}&m_{I}^{2}\\[2.0pt] m_{I}^{2}&-p^{2}\end{bmatrix}\begin{pmatrix}\vec{\phi}_{1}(-p)\\ \vec{\phi}_{2}(-p)\end{pmatrix}$ (57)
$+\dots\,.$
@@
IX Dualities among $SO(N)$ Chern–Simons theories from many Majoranas
Consider now the bosonization of an arbitrary number of Majorana flavours. The partition function is given by
$Z_{N_{f}}=\int\prod_{i=1}^{N_{f}}\mathcal{D}\xi_{i}\,\mathcal{D}\bar{\xi}_{i}\;\exp\!\left[i\int d^{3}x\,\sum_{i=1}^{N_{f}}i\bar{\xi}_{i}\cancel{\partial}\xi_{i}\right].$ (61)
which is consistent with
$Z_{N_{f}}=(-)^{s(N_{f})}\,\underbrace{Z_{1}\cdots Z_{1}}_{N_{f}~\text{times}}\,$, (62)
where $s(N_{f})$ counts the number of anticommutations needed to reorder the Grassmann measures i.e.
$s(N_{f})=\sum_{n=1}^{N_{f}}(N_{f}-n)=\frac{N_{f}(N_{f}-1)}{2}\,.$ (63)
Physical observables are insensitive to the overall sign in (62), so a theory of $N_{f}$ free Majoranas may be treated as $N_{f}$ copies of the single-flavour theory. This leads to the following relation between the two bosons side of the duality:
$\sum_{i=1}^{N_{f}}|D_{a}\vec{\phi}_{i}|^{2}-\tilde{V}(\vec{\phi}_{1},\ldots,\vec{\phi}_{N_{f}})+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}$
$\longleftrightarrow\sum_{i=1}^{N_{f}}\!\Big(|D_{a_{i}}z_{i}|^{2}-|z_{i}|^{4}+\text{CS}_{SO(N_{i})}[a_{i}]_{1}+N_{i}\text{CS}_{g}\Big)\,.$
In principle the ranks $N_{i}$ are unrelated. Combining with the constraints (14) yields
$N\geq N_{f}+2\,,\qquad N_{i}\geq 3\,.$ (64)
For simplicity, take $N_{i}=\bar{N}$} for all $i$. The conjectured duality becomes
$SO(N)_{1}\;+\;N_{f}~\text{coupled real scalars}\;+\;N\text{CS}_{g}$
$\;\longleftrightarrow\;$ $N_{f}\!\left(SO(\bar{N})_{1}\;+\;\text{one Wilson--Fisher scalar}\right)$
$\;+\;N_{f}\bar{N}\,\text{CS}_{g}\,.$ (65)
To chart the phases we again introduce an auxiliary scalar $\varphi$ and identify the relevant deformation as
$\sum_{i=1}^{N_{f}}\vec{\phi}_{i}^{2}\;\longleftrightarrow\;\sum_{i=1}^{N_{f}}|z_{i}|^{2}\,.$ (66)
Near the critical point,
$\sum_{i=1}^{N_{f}}\!\Big(|D_{a}\vec{\phi}_{i}|^{2}-h\varphi\,\vec{\phi}_{i}^{2}\Big)-\tilde{V}(\vec{\phi}_{1},\ldots,\vec{\phi}_{N_{f}})$
$+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}-v(\varphi)$
$\longleftrightarrow\sum_{i=1}^{N_{f}}\!\Big(|D_{a_{i}}z_{i}|^{2}-h\varphi\,|z_{i}|^{2}-|z_{i}|^{4}+\text{CS}_{SO(\bar{N})}[a_{i}]_{1}\Big)$
$+\;N_{f}\bar{N}\,\text{CS}_{g}-v(\varphi)\,$, (67)
with h>0{h>0}. Promoting $\varphi$ to a dynamical field and focusing on the phase with $\varphi_{0}>0$}, all scalars are massive and can be integrated out, leaving
$SO(N)_{1}+N\text{CS}_{g}\;\longleftrightarrow\;N_{f}\,SO(\bar{N})_{1}+N_{f}\bar{N}\,\text{CS}_{g}\,.$
Both bulk theories are fully gapped; boundary chiral modes are cancelled by the gravitational counterterms.
For $\varphi_{0}<0$}, the $\vec{\phi}_{i}$ condense and Higgs $SO(N)\!\to\!SO(N-N_{f})$; on the right, each $z_{i}$ condenses, breaking $SO(\bar{N})\!\to\!SO(\bar{N}-1)$ in every copy. The IR theories are
$SO(N-N_{f})_{1}+N\text{CS}_{g}\;\longleftrightarrow\;N_{f}\,SO(\bar{N}-1)_{1}+N_{f}\bar{N}\,\text{CS}_{g}\,.$
Both are again gapped in the bulk, while the boundary carries $c_{-}=\tfrac{N_{f}}{2}$}. Thus (67) interpolates between a phase with vanishing thermal Hall response and one with $k_{xy}=\tfrac{N_{f}\pi}{12}T$} at low temperature. Dropping the gravitational terms for brevity, we summarize the IR dualities as
$SO(N)_{1}$ $\;\longleftrightarrow\;N_{f}\,SO(\bar{N})_{1}\,$, (68)
$SO(N-N_{f})_{1}$ $\;\longleftrightarrow\;N_{f}\,SO(\bar{N}-1)_{1}\,$, (69)
valid for $N\geq N_{f}+2$ and $\bar{N}\geq 3$.
X Conclusions
@@
In this work we have extended the operator-identification method of Ref. [Witczak-Krempa:2023lgm], originally developed for Dirac fermions, to the framework of Majorana bosonization. Starting from the decomposition of a Dirac spinor into two neutral components, we analyzed how this structure translates within the $SO(N)$ Chern–Simons formulation of Ref. [Metlitski:2016dht, Aharony:2016jvv]. The correspondence between the two bosonizations naturally leads to an orthogonal (IR) description of the Chern–Simons–Higgs model, in which boundary conditions for Majorana flavours play a central role. In particular, chirality-preserving boundary conditions map to specific constraints on monopole operators: when scalar fields satisfy these relations, monopoles may exist at the boundary, whereas anomaly matching requires their absence otherwise.
@@
Building on this framework, we identified new critical points and phase transitions. In Section VII we described an orthogonal analogue of the transition between a trivial insulator and a quantum spin Hall phase, while in Section VIII we analyzed a pairing-type deformation of the Dirac seed duality that produces a VEV-dependent Higgs phase. A distinctive feature of all these results is the flexibility in the rank of the gauge group appearing in Majorana bosonization. This freedom was exploited in Section IX, where an arbitrary number of Majorana fermions was bosonized in two inequivalent ways, leading to the dualities in Eqs. (68)–(69). Despite the apparent triviality of level-one $SO(N)$ Chern–Simons theories in the bulk, the corresponding Wess–Zumino–Witten edge theories may encode non-trivial dynamics and deserve a dedicated analysis.
@@
Our study also highlights the limitations of the operator-identification method: although it efficiently reveals new candidate fixed points, it does not by itself guarantee that the associated renormalization-group flows coincide. A more complete RG treatment could refine or constrain the parameter range in Eq. (64) and clarify possible mismatches between the bosonic and fermionic trajectories. Likewise, the consistency of reusing the $N_{f}=2$ bosonization twice in the Ising–Gross–Neveu construction remains an open technical question.
@@
Future developments should focus on two main directions. First, the ambiguity in the rank of the gauge group within Majorana bosonization should be systematically investigated, possibly through explicit RG analyses or lattice realizations. Second, the framework could be expanded to include further critical points beyond the Dirac–Majorana connection. Exploring known topological phases supporting Majorana modes, such as Kitaev spin liquids [Kitaev:2005hzj], or identifying new classes of IR dualities among orthogonal Chern–Simons–matter theories may provide deeper insight into the landscape of three-dimensional topological quantum field theories.

### v3

@@ 8 Pairing transition
where $\psi^{2}\equiv\epsilon^{\alpha\beta}\psi_{\alpha}\psi_{\beta}$} with $\epsilon^{12}=-\epsilon^{21}=1$. On the bosonic side, using the identification of the Dirac operator with a monopole operator [12], the quadratic operator maps as
@@
$\mathcal{L}_{z}=|D_{a}z|^{2}-|z|^{4}+\frac{1}{4\pi}a\wedge da+\frac{1}{2\pi}A\wedge da\;-\big(h\varphi\,\mathcal{M}_{2}+\text{c.c.}\big)-v(\varphi)\,.$ (39)
@@
Since $\mathscr{M}_{a}$ is unitary [23],
@@
$\mathcal{L}_{\phi}=|D_{a}\vec{\phi}_{1}|^{2}+|D_{a}\vec{\phi}_{2}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})-h\Big[\varphi_{R}\big(\vec{\phi}_{1}^{2}-\vec{\phi}_{2}^{2}\big)-2\varphi_{I}\vec{\phi}_{1}\cdot\vec{\phi}_{2}\Big]\\ -v(\varphi)+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}\,.$ (51)
@@
$\mathcal{L}_{\phi}=|D_{a}\vec{\phi}_{1}|^{2}+|D_{a}\vec{\phi}_{2}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}-h\varphi_{R,0}\big(\vec{\phi}_{1}^{2}-\vec{\phi}_{2}^{2}\big)\,.$ (53)
@@
$\mathcal{L}_{\phi}=|D_{a}\vec{\phi}_{1}|^{2}+|D_{a}\vec{\phi}_{2}|^{2}-\tilde{V}(\vec{\phi}_{1},\vec{\phi}_{2})+\text{CS}_{SO(N)}[a]_{1}+N\text{CS}_{g}+2h\varphi_{I,0}\vec{\phi}_{1}\cdot\vec{\phi}_{2}\,.$ (55)
@@
$S_{\phi}=\int d^{3}x\,\Big[\partial_{\mu}\vec{\phi}_{1}\cdot\partial^{\mu}\vec{\phi}_{1}+\partial_{\mu}\vec{\phi}_{2}\cdot\partial^{\mu}\vec{\phi}_{2}+2h\varphi_{I,0}\vec{\phi}_{1}\cdot\vec{\phi}_{2}+\dots\Big]\,.$ (56)
@@
$S_{\phi}=\int\frac{d^{3}p}{(2\pi)^{3}}\begin{pmatrix}\vec{\phi}_{1}^{\mathrm{T}}(p)&\vec{\phi}_{2}^{\mathrm{T}}(p)\end{pmatrix}\begin{bmatrix}-p^{2}&m_{I}^{2}\\[2.0pt] m_{I}^{2}&-p^{2}\end{bmatrix}\begin{pmatrix}\vec{\phi}_{1}(-p)\\ \vec{\phi}_{2}(-p)\end{pmatrix}\;$$\dots\,$. (57)
@@
9 Conclusions
@@
In this work we have extended the operator-identification method of Ref. [24], originally developed for Dirac fermions, to the framework of Majorana bosonization. Starting from the decomposition of a Dirac spinor into two neutral components, we analyzed how this structure translates within the $SO(N)$ Chern–Simons formulation of Ref. [14, 1]. The correspondence between the two bosonizations naturally leads to an orthogonal (IR) description of the Chern–Simons–Higgs model, in which boundary conditions for Majorana flavours play a central role. In particular, chirality-preserving boundary conditions map to specific constraints on monopole operators: when scalar fields satisfy these relations, monopoles may exist at the boundary, whereas anomaly matching requires their absence otherwise.
@@
Building on this framework, we identified new critical points and phase transitions. In Section 7 we described an orthogonal analogue of the transition between a trivial insulator and a quantum spin Hall phase, while in Section 8 we analyzed a pairing-type deformation of the Dirac seed duality that produces a VEV-dependent Higgs phase. A distinctive feature of all these results is the flexibility in the rank of the gauge group appearing in Majorana bosonization.
@@
Our study also highlights the limitations of the operator-identification method: although it efficiently reveals new candidate fixed points, it does not by itself guarantee that the associated renormalization-group flows coincide. By considering a free theory of multiple Majorana flavours, indeed, problematic results regarding the flexibility in the rank of the gauge group and the number of bosonized fermions may appear. A more complete RG treatment could refine or constrain the parameter range and clarify possible mismatches between the bosonic and fermionic trajectories. Likewise, the consistency of reusing the $N_{f}=2$ bosonization twice in the Ising–Gross–Neveu construction remains an open technical question.
@@
Future developments should focus on two main directions. First, the ambiguity in the rank of the gauge group within Majorana bosonization should be systematically investigated, possibly through explicit RG analyses or lattice realizations. Second, the framework could be expanded to include further critical points beyond the Dirac–Majorana connection. Exploring known topological phases supporting Majorana modes, such as Kitaev spin liquids [13], or identifying new classes of IR dualities among orthogonal Chern–Simons–matter theories may provide deeper insight into the landscape of three-dimensional topological quantum field theories.